# Twisted compactifications and conformal defects from ISO(3)$\times$×U(1) $F(4)$ gauged supergravity

### v1

@@ 3.2 Compactifications to two dimensions
We recall that this $SO(3)$ subgroup is generated by $X_{r}$. There are two $SO(3)$ singlet scalars parametrized by the following coset representative, see more detail in [50],
@@
It should be noted that only two of these projectors are independent, so the $AdS_{3}\times\mathcal{M}_{3}$ solutions will preserve four supercharges. This unbroken supersymmetry would correspond to $N=(1,1)$ superconformal symmetry in two dimensions, see also the discussion in [29].
@@
It can also be checked that all these equations satisfy the field equations. It turns out that these equations admit only one $AdS_{3}\times\mathcal{M}_{3}$ fixed point given by
@@
As in the previous case, the fixed point exists only for $\kappa=-1$ giving rise to an $AdS_{3}\times H^{3}$ solution. Examples of numerical solutions interpolating between this geometry and the $AdS_{6}$ vacuum are given in figure 3 for $g=3m$ and $m=\frac{1}{4}$ (red), $m=\frac{1}{2}$ (green), $m=1$ (blue). Since $\varphi=0$ both at the $AdS_{3}\times H^{3}$ fixed point and along the flow, the solutions can also be regarded as solutions of $ISO(3)$ $F(4)$ gauged supergravity coupled to three vector multiplets. The solutions describe RG flows from the five-dimensional $N=2$ SCFT to a two-dimensional $N=(1,1)$ SCFT in the IR. Equivalently, the solutions can also be interpreted as supersymmetric black strings in asymptotically $AdS_{6}$ space.
As in the previous case, $\sigma$ participate in the flows with the behavior near the $AdS_{6}$ vacuum
@@
Figure 3: Supersymmetric RG flows from $N=2$ SCFT in five dimensions to $N=(1,1)$ SCFT in two dimensions from $ISO(3)$ $F(4)$ gauged supergravity.
@@ 4.1 Line defects
In this equation, $\epsilon_{0}$ is a two-component constant spinor while $\eta$ and $\chi^{\pm}$ are respectively Killing spinors on $AdS_{2}$ and $S^{3}$ satisfying
@@
Following [43], we use the following form of gamma matrices
@@
$b=-\frac{e^{f+\sigma}}{\ell m}\quad\textrm{and}\qquad R=\frac{2\ell e^{2f-h+\sigma}}{2e^{f+\sigma}-\ell mb}\,.$ (87)
It is useful to point out that these two constraints result from the fact that $\delta\chi_{A}$, $\delta\psi_{Aa}$ and $\delta\psi_{Ai}$ lead to three different BPS equations for $b(r)$. Taking one of them as a flow equation for $b(r)$ and requiring this equation be equivalent to the remaining two equations give rise to the aforementioned two algebraic constraints, see more detail in [43]. In the BPS equations, we have also chosen a particular sign choice to identify the $AdS_{6}$ vacuum as $r\rightarrow\infty$.
Since the BPS equations for $h^{\prime}$ and $f^{\prime}$ are the same, we imediately find that $f=h+C$ for a constant $C$. The constant $C$ can be chosen to be zero by requiring that the solution becomes locally $AdS_{6}$ geometry with $f\sim h\sim\frac{r}{L}$. Therefore, we will set $f(r)=h(r)$ from now on. We have also verified that the BPS equations are compatible with the two algebraic constraints and the corresponding second-ordered field equations. We are not able to analytically solve these equations in full generality. This requires some numerical analysis which can always be achived by suitable boundary conditions. We will instead consider some interesting solutions that can be obtained analytically within particular subtruncations. These solutions could be much more useful in most proposes.
@@
$b$ $=$ $-\frac{(\sigma_{0}+3m\coth\phi_{5})^{\frac{1}{3}}}{Lmg^{\frac{1}{4}}},\qquad Y=e^{\frac{f}{2}}$ (88)
with $\sigma_{0}$ being an integration constant and $\rho$ defined by $\frac{d\rho}{dr}=e^{\sigma}$. First of all, we notice that the solution approaches the $AdS_{6}$ vacuum with $\phi_{5}=0$ and $\sigma=\frac{1}{4}\ln\left[\frac{3m}{g}\right]$ for any value of $\sigma_{0}$. This solution takes the same form as the holographic RG flow studied in [50] with the presence of a non-vanishing two-form field. This type of solutions is called charged domain walls in [37] and [38]. Near the $AdS_{6}$ vacuum, we find
@@
$\sigma\sim\frac{1}{4}\ln(\rho-\rho_{0}),\qquad f\sim\frac{5}{12}\ln(\rho-\rho_{0}),\qquad b\sim-\frac{6^{\frac{1}{3}}(\rho-\rho_{0})^{\frac{2}{3}}}{Lg^{\frac{1}{4}}m^{\frac{2}{3}}}$ (91)
@@
$\sigma\sim-\frac{1}{4}\ln(\rho-\rho_{0}),\qquad f\sim\frac{1}{4}\ln(\rho-\rho_{0}),\qquad b\sim-\frac{(3m+\sigma_{0})^{\frac{1}{3}}}{Lg^{\frac{1}{4}}m}$ (92)
@@
$b$ $=$ $-\frac{e^{-\frac{4}{3}\phi_{1}}(3m-\tilde{\sigma}_{0}e^{4\phi_{1}})^{\frac{1}{3}}}{Lm(2g)^{\frac{1}{4}}},\qquad Y=e^{\frac{f}{2}}$ (93)
@@
$f\sim\frac{1}{12}\ln(\rho-\rho_{0}),\qquad b\sim-\frac{3^{\frac{1}{3}}}{(2g)^{\frac{1}{4}}m^{\frac{2}{3}}L}\,.$ (95)
@@ 4.2 Surface defects
The scalar fields are given by the coset representative given in (27).
@@
As in the case of line defects, we will consider two types of solutions that can be analytically obtained in particular subtruncations. The first type of solutions is found by setting $\phi_{2}=\phi_{1}=0$ which leads to the following solution
@@
$b$ $=$ $-\frac{2}{Lmg^{\frac{1}{4}}}(\sigma_{0}+3m\coth\phi_{5})^{\frac{1}{3}},\qquad Y=e^{\frac{f}{2}}$ (111)
@@
$b$ $=$ $-\frac{2^{\frac{3}{4}}e^{-\frac{4}{3}\phi_{1}}(3m-\tilde{\sigma}_{0}e^{4\phi_{1}})^{\frac{1}{3}}}{Lmg^{\frac{1}{4}}},\qquad Y=e^{\frac{f}{2}}\,.$ (112)
with the same coordinate $\rho$ as in the previous solution. All of these solutions take the same form as the solutions for line defects considered above with the only difference being a numerical factor in the solution for $b$. Asymptotic behaviors are the same as in the case of line defects, and we will not repeat these again. We simply point out that the surface defects also arise from turning on a dimension-$4$ operator dual to the two-form field as in the case of line defects. Although the solutions in both cases take the same form, we expect the two types of solutions lead to different uplifted ten-dimensional solutions. In particular, in the case of line defects, we have an electric component of the two-form field while the magnetic component appears in the case of surface defects.
@@ 5 Conclusions
We have studied a number of supersymmetric solutions from matter-coupled $F(4)$ gauged supergravity with $ISO(3)\times U(1)$ gauge group. The first class of these solutions takes the form of RG flow across dimensions and holographically describes twisted compactifications of the five-dimensional $N=2$ SCFT on $H^{2}$ and $H^{3}$ to three- and two-dimensional SCFTs in the IR. These solutions have been obtained by performing standard topological twists using $SO(2)\times U(1)$ and $SO(3)$ gauge fields. Since the $ISO(3)\times U(1)$ $F(4)$ gauged supergravity is a consistent truncation of type IIB theory on $S^{2}\times\Sigma$, the $AdS_{4}\times H^{2}$ and $AdS_{3}\times H^{3}$ solutions found in this paper lead to new supersymmetric $AdS_{4}\times H^{2}\times S^{2}\times\Sigma$ and $AdS_{3}\times H^{3}\times S^{2}\times\Sigma$ solutions in type IIB theory.
The second class of solutions obtained in this paper takes the form of “charged” curved domain walls with $AdS_{2}\times S^{3}$ and $AdS_{3}\times S^{2}$ slicing. The solutions have been found by turning on the two-form field in the gravity multiplet to support the curvature of the domain walls. The solutions have a holographic interpretation as conformal line and surface defects within the $N=2$ SCFT in five dimensions. These defects arise from a source term for a dimension-$4$ operator dual to the two-form field. The solutions also have a very similar structure as those found in compact gauge groups studied in [43]. In particular, the solutions are mainly holographic RG flows in the presence of a running two-form field.
Although the matter-coupled $ISO(3)\times U(1)$ $F(4)$ gauged supergravity considered here can be embedded in type IIB theory, the complete truncation ansatz has not been constructed to date. It would be interesting to work this out using the formalism of exceptional field theories and use the resulting truncation ansatz to uplift the solutions found in this paper to ten dimensions. This could lead to new $AdS_{4}\times H^{2}$ and $AdS_{3}\times H^{3}$ solutions in type IIB theory as mentioned above as well as possible brane configurations whose near horizon geometries given by these solutions. It would also be interesting to identify the dual $N=2$ SCFT in five dimensions together with field theory descriptions of conformal line and surface defects dual to the solutions found here in terms of position-dependent deformations.

### v3

@@ 3.2 Compactifications to two dimensions
In this case, $dx^{2}_{1,1}$ is the flat metric on two-dimensional Minkowski space with the signature (+)(-+)( - + ).
@@
We recall that this $SO(3)$ subgroup is generated by $X_{r}$ for $r=1,2,3$. There are two $SO(3)$ singlet scalars parametrized by the following coset representative, see more detail in [50],
@@
With all these conditions together with the ansatz for the $SO(3)$ gauge fields given in (51), all the terms appearing in (55) cancel among each other.
It should be noted that only two of the projectors in (57) are independent, so the $AdS_{3}\times\mathcal{M}_{3}$ solutions will preserve four supercharges. In the dual five-dimensional $N=2$ SCFT, the $16$ supercharges transform under $SO(1,4)\times SO(3)_{R}$ as $(\mathbf{4},\mathbf{2})$ and $(\bar{\mathbf{4}},\mathbf{2})$. Under the decomposition to two dimensions, we have $SO(1,4)\rightarrow SO(1,1)\times SO(3)$ under which $\mathbf{4}\rightarrow\mathbf{2}_{1}+\mathbf{2}_{-1}$. The original $16$ supercharges then transform under $SO(1,1)\times SO(3)\times SO(3)_{R}$ as
$(\mathbf{4},\mathbf{2})+(\bar{\mathbf{4}},\mathbf{2})\rightarrow 2\times(\mathbf{2}_{1}+\mathbf{2}_{-1},\mathbf{2}).$ (58)
After imposing the topological twist by identifying the $SO(3)$ with the $SO(3)_{R}$ symmetry, we end up with the following decomposition
$(\mathbf{4},\mathbf{2})+(\bar{\mathbf{4}},\mathbf{2})\rightarrow 2\times(\mathbf{1}_{1}+\mathbf{1}_{-1}+\mathbf{3}_{1}+\mathbf{3}_{-1})$ (59)
under $SO(1,1)\times SO(3)_{\textrm{diag}}$. The unbroken supersymmetry is then generated by the singlet supercharges and corresponds to $N=(1,1)$ superconformal symmetry in two dimensions, see also the discussion in [26].
@@
It can also be checked that all these equations satisfy the field equations. As in the previous case, for $a=0$, these equations reduce to those considered in [50].
It turns out that these equations admit only one $AdS_{3}\times\mathcal{M}_{3}$ fixed point given by
@@
The fixed point exists only for $\kappa=-1$ giving rise to an $AdS_{3}\times H^{3}$ solution. Examples of numerical solutions interpolating between this geometry and the $AdS_{6}$ vacuum are given in figure 3 for $g=3m$ and $m=\frac{1}{4}$ (red), $m=\frac{1}{2}$ (green), $m=1$ (blue). Since $\varphi=0$ both at the $AdS_{3}\times H^{3}$ fixed point and along the flow, the solutions can also be regarded as solutions of $ISO(3)$ $F(4)$ gauged supergravity coupled to three vector multiplets. The solutions describe RG flows from the five-dimensional $N=2$ SCFT to a two-dimensional $N=(1,1)$ SCFT in the IR. Equivalently, the solutions can also be interpreted as supersymmetric black strings in asymptotically $AdS_{6}$ space.
As in the previous case, $\sigma$ participates in the flows with the behavior near the $AdS_{6}$ vacuum
@@
Figure 3: Supersymmetric RG flows from $N=2$ SCFT in five dimensions (dual to the $AdS_{6}$ vacuum (21)) to $N=(1,1)$ SCFT in two dimensions (dual to the $AdS_{3}\times H^{3}$ vacuum (66)) from $ISO(3)$ $F(4)$ gauged supergravity with $g=3m$ and $m=\frac{1}{4}$ (red), $m=\frac{1}{2}$ (green), $m=1$ (blue).
@@ 4.1 Line defects
In this equation, $\epsilon_{0}$ is a two-component constant spinor while $\chi^{\pm}$ and $\eta$ are respectively Killing spinors on $AdS_{2}$ and $S^{3}$ satisfying
@@
As in [43], we use the following form of gamma matrices
@@
$b=-\frac{e^{f+\sigma}}{\ell m}\qquad\textrm{and}\qquad R=\frac{2\ell e^{2f-h+\sigma}}{2e^{f+\sigma}-\ell mb}\,.$ (89)
In all the BPS equations given above, we have replaced the constant $R$ by using the second constraint given in (89). It is also useful to point out that the two algebraic constraints result from the fact that $\delta\chi_{A}$, $\delta\psi_{Aa}$ and $\delta\psi_{Ai}$ lead to three different BPS equations for $b(r)$. Taking one of them as a flow equation for $b(r)$ and requiring this equation be equivalent to the remaining two equations give rise to the aforementioned two algebraic constraints, see more detail in [43]. In the BPS equations, we have also chosen a particular sign choice to identify the $AdS_{6}$ vacuum as $r\rightarrow\infty$.
Since the BPS equations for $h^{\prime}$ and $f^{\prime}$ are the same, we immediately find that $f=h+C$ for a constant $C$. The constant $C$ can be chosen to be zero by requiring that the solution becomes locally $AdS_{6}$ geometry with $f\sim h\sim\frac{r}{L}$. Therefore, we will set $f(r)=h(r)$ from now on. We have also verified that the BPS equations are compatible with the two algebraic constraints and the corresponding second-ordered field equations. We are not able to analytically solve these equations in full generality. This requires some numerical analysis which can always be achieved by suitable boundary conditions. We will instead consider some interesting solutions that can be obtained analytically within particular subtruncations. These solutions could be much more useful in most purposes.
@@
$b$ $=$ $-\frac{(\sigma_{0}+3m\coth\phi_{5})^{\frac{1}{3}}}{\ell mg^{\frac{1}{4}}},\qquad Y=e^{\frac{f}{2}}$ (90)
with $\sigma_{0}$ being an integration constant and $\rho$ defined by $\frac{d\rho}{dr}=e^{\sigma}$.
First of all, we notice that the solution approaches the $AdS_{6}$ vacuum with $\phi_{5}=0$ and $\sigma=\frac{1}{4}\ln\left[\frac{3m}{g}\right]$ for any value of $\sigma_{0}$. This solution takes the same form as the holographic RG flow studied in [50] with a non-vanishing two-form field. This type of solutions is called charged domain walls in [34] and [35]. Near the $AdS_{6}$ vacuum, we find
@@
$\sigma\sim\frac{1}{4}\ln(\rho-\rho_{0}),\qquad f\sim\frac{5}{12}\ln(\rho-\rho_{0}),\qquad b\sim-\frac{6^{\frac{1}{3}}(\rho-\rho_{0})^{\frac{2}{3}}}{\ell g^{\frac{1}{4}}m^{\frac{2}{3}}}$ (93)
@@
$\sigma\sim-\frac{1}{4}\ln(\rho-\rho_{0}),\qquad f\sim\frac{1}{4}\ln(\rho-\rho_{0}),\qquad b\sim-\frac{(3m+\sigma_{0})^{\frac{1}{3}}}{\ell g^{\frac{1}{4}}m}$ (94)
@@
$b$ $=$ $-\frac{e^{-\frac{4}{3}\phi_{1}}(3m-\tilde{\sigma}_{0}e^{4\phi_{1}})^{\frac{1}{3}}}{\ell m(2g)^{\frac{1}{4}}},\qquad Y=e^{\frac{f}{2}}$ (95)
@@
$f\sim\frac{1}{12}\ln(\rho-\rho_{0}),\qquad b\sim-\frac{3^{\frac{1}{3}}}{(2g)^{\frac{1}{4}}m^{\frac{2}{3}}\ell}\,.$ (97)
@@ 4.2 Surface defects
The scalar fields are given by the coset representative in (27).
@@
As in the case of line defects, we have replaced the constant $R$ in the BPS equations by using the second constraint given in (112). We now consider two types of solutions that can be analytically obtained in particular subtruncations. The first type of solutions is found by setting $\phi_{2}=\phi_{1}=0$ which leads to the following solution
@@
$b$ $=$ $-\frac{2}{\ell mg^{\frac{1}{4}}}(\sigma_{0}+3m\coth\phi_{5})^{\frac{1}{3}},\qquad Y=e^{\frac{f}{2}}$ (113)
@@
$b$ $=$ $-\frac{2^{\frac{3}{4}}e^{-\frac{4}{3}\phi_{1}}(3m-\tilde{\sigma}_{0}e^{4\phi_{1}})^{\frac{1}{3}}}{\ell mg^{\frac{1}{4}}},\qquad Y=e^{\frac{f}{2}}\,.$ (114)
with the same coordinate $\rho$ as in the previous solution. The integration constant $\tilde{\sigma}_{0}$ needs to be chosen as $\tilde{\sigma}_{0}=3m$ in order for the solution to approach the $AdS_{6}$ vacuum. All of these solutions take the same form as the solutions for line defects considered above with the only difference being a numerical factor in the solution for $b$. Asymptotic behaviors are the same as in the case of line defects, and we will not repeat these again. We simply point out that the surface defects also arise from turning on a dimension-$4$ operator dual to the two-form field as in the case of line defects. Although the solutions in both cases take the same form, we expect that the two types of solutions lead to different uplifted ten-dimensional solutions. In particular, in the case of line defects, we have an electric component of the two-form field while the magnetic component appears in the case of surface defects.
We end this section by giving some comments regarding the solutions of line and surface defects in matter-coupled $F(4)$ gauged supergravity with a compact $SU(2)\times U(1)$ gauge group considered in [43] and a non-semisimple $ISO(3)\times U(1)$ gauge group studied in this paper. We note that although the $SO(2)\times U(1)$ symmetric solutions given here take the same from as those in [43], the solutions presented in this paper are obtained within a particular subtruncation with $\phi_{1}=\phi_{2}=0$. The more general defect solutions, obtained by including these two scalars dual to irrelevant operators in the five-dimensional $N=2$ SCFT, are certainly different from the solutions found in [43]. The latter are, on the other hand, the most general solutions within the $SO(2)\times U(1)$ symmetric sector not in any particular subtruncations. Therefore, the crucial difference between the solutions from compact and non-semisimple gauge groups lies in the presence of the dual irrelevant operators.
@@ 5 Conclusions
We have studied a number of supersymmetric solutions from matter-coupled $F(4)$ gauged supergravity with $ISO(3)\times U(1)$ gauge group. The first class of these solutions takes the form of RG flows across dimensions and holographically describes twisted compactifications of the five-dimensional $N=2$ SCFT on $H^{2}$ and $H^{3}$ to three- and two-dimensional SCFTs in the IR. These solutions have been obtained by performing standard topological twists using $SO(2)\times U(1)$ and $SO(3)$ gauge fields. Since the $ISO(3)\times U(1)$ $F(4)$ gauged supergravity is a consistent truncation of type IIB theory on $S^{2}\times\Sigma$, the $AdS_{4}\times H^{2}$ and $AdS_{3}\times H^{3}$ solutions found in this paper could lead to new supersymmetric $AdS_{4}\times H^{2}\times S^{2}\times\Sigma$ and $AdS_{3}\times H^{3}\times S^{2}\times\Sigma$ solutions in type IIB theory.
The second class of solutions obtained in this paper takes the form of “charged” curved domain walls with $AdS_{2}\times S^{3}$ and $AdS_{3}\times S^{2}$ slicing. These solutions have been found by turning on the two-form field in the gravity multiplet to support the curvature of the domain walls. The solutions have a holographic interpretation as conformal line and surface defects within the $N=2$ SCFT in five dimensions. These defects arise from a source term for a dimension-$4$ operator dual to the two-form field. The solutions also have a very similar structure as those found in compact gauge groups studied in [43]. In particular, the solutions are mainly holographic RG flows in the presence of a running two-form field.
Although the matter-coupled $ISO(3)\times U(1)$ $F(4)$ gauged supergravity considered here can be embedded in type IIB theory, the complete truncation ansatz has not been constructed to date. It would be interesting to work this out using the formalism of exceptional field theories and use the resulting truncation ansatz to uplift the solutions found in this paper to ten dimensions. This could lead to new $AdS_{4}\times H^{2}$ and $AdS_{3}\times H^{3}$ solutions in type IIB theory as mentioned above as well as possible brane configurations with near horizon geometries given by these solutions. It would also be interesting to identify the dual $N=2$ SCFT in five dimensions together with field theory descriptions of conformal line and surface defects dual to the solutions found here in terms of position-dependent deformations.